OpenStudy (anonymous):

The price of a community pool membership has a one-time sign-up fee and a monthly fee. The price can be modeled by the function y = 20x + 50 where x is the number of months. What is the slope and what does it represent? A. 20; it represents the monthly fee B. 20; it represents the one-time sign-up fee C. 50; it represents the monthly fee D. 50; it represents the one-time sign-up fee

OpenStudy (jusaquikie):

ok think of y as the total amount you spemd on the pool, which would change with time and which would be a one time cost?

OpenStudy (jusaquikie):

now the slope would be the part that changes with time, the other would just be where what the cost would be at time =0 or the y intercept or the sign-up fee, hint hint

OpenStudy (anonymous):

When you first get to the pool, your number of months there is 0 (x=0). So y = 20x + 50 = 20(0) + 50 = 50. So that means 50 must be the sign-up fee. Also, the first month (x=1) you've payed y = 20(1)+ 50 = 70. That means you've payed 20 more than before. So that 20 must be the monthly fee.